OpenStudy (anonymous):

The solution to square root of the quantity 2 times x plus 3 minus square root of the quantity x plus 1 equals 1 is x = 3. True or false

OpenStudy (whpalmer4):

Well, plug x=3 into the equation. Does it work?

OpenStudy (whpalmer4):

\[\sqrt{2x+3} -\sqrt{x+1} = 1\]\[\sqrt{2*3+3}-\sqrt{3+1} = 1\] Does it work? Now, your instructor may be trying to set a trap for you here.

OpenStudy (anonymous):

Yes it does! Can you keep helping me?

OpenStudy (anonymous):

Solve square root of 3 times x equals square root of the quantity 4 times x minus 1.

OpenStudy (whpalmer4):

The problem says "The solution is..." which implies that there is only 1 solution. Actually, there's another solution, so if you take the problem strictly at its word, the answer would be False, because that is not the only solution. Try out \(x=-1\): \[\sqrt{2(-1)+3} - \sqrt{-1+1} = 1\]\[\sqrt{-2+3} - \sqrt{0} = 1\]\[\sqrt{1} = 1\]\[1 = 1\] It turns out that this equation has two solutions, not just one. However, if you haven't covered having multiple solutions, it may just be a case of sloppy problem writing.

OpenStudy (whpalmer4):

\[\sqrt{3x} = \sqrt{4x-1}\]Is that the new problem?

OpenStudy (whpalmer4):

What do you get if you square both sides of the equation?

OpenStudy (anonymous):

Idk what you mean?

OpenStudy (whpalmer4):

You don't know what I mean by what, precisely? Squaring both sides of the equation?

OpenStudy (anonymous):

yes what

OpenStudy (anonymous):

does that mean

OpenStudy (anonymous):

??

OpenStudy (whpalmer4):

Multiply each side by itself \[\sqrt{3x}*\sqrt{3x} = \sqrt{4x-1}*\sqrt{4x-1}\]Can you simplify that?

OpenStudy (anonymous):

Ugh no

OpenStudy (anonymous):

\[\sqrt{9x ^{2}}\] is that it for the first part?

OpenStudy (whpalmer4):

\[\sqrt{u}*\sqrt{u} = \sqrt{u*u} = \sqrt{u^2}\]If \(u>0\), \[\sqrt{u^2} =u\]So assuming that \(x>0\), we can just write \[\sqrt{3x}*\sqrt{3x} = \sqrt{4x-1}*\sqrt{4x-1}\]\[3x = 4x-1\]And I hope you can solve that for the value of \(x\)

OpenStudy (anonymous):

is it -2

OpenStudy (whpalmer4):

:-( Does \(3(-2) = 4(-2) -1\)?

OpenStudy (anonymous):

-1

OpenStudy (whpalmer4):

Again, try your solution before asking me if you are correct. Does \[3(-1) = 4(-1) - 1\]

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

That's better. Now, because we squared a square root while solving, we need to test our solution in the original equation. It may or may not work! If it does not, it is what is called an extraneous solution. \[\sqrt{3(1)} = \sqrt{4(1) -1}\]Is that true?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Solve square root of the quantity two x plus one equals x plus one for x.

OpenStudy (whpalmer4):

\[\sqrt{2x+1} = \sqrt{x+1}\]Okay, what's your first step here going to be?

OpenStudy (whpalmer4):

(you ought to be able to give me the answer to this one with a few seconds thought, no algebra needed)

OpenStudy (anonymous):

I think it's 0

OpenStudy (whpalmer4):

Yes, that is correct.

OpenStudy (anonymous):

Solve five equals square root of the quantity x squared plus sixteen for x.

OpenStudy (whpalmer4):

But how would you go through the motions of solving it?

OpenStudy (anonymous):

I think the answer is 3

OpenStudy (whpalmer4):

\[5 = \sqrt{x^2+16}\] Okay, yes, x = 3 is a solution. Is it the only solution?

OpenStudy (anonymous):

\[\pm3\]

OpenStudy (whpalmer4):

Much better.

OpenStudy (anonymous):

Solve square root of the quantity k minus seven equals the square root of k, minus one for k.

OpenStudy (whpalmer4):

Okay, it's time for you to learn to use the Equation editor button...

OpenStudy (anonymous):

scary

OpenStudy (whpalmer4):

\[\sqrt{k-7} = \sqrt{k}-1\]

OpenStudy (whpalmer4):

How are you going to do this?

OpenStudy (anonymous):

I think it's 16

OpenStudy (whpalmer4):

Okay.

OpenStudy (anonymous):

Is that the answer?

OpenStudy (whpalmer4):

Does it work?

OpenStudy (anonymous):

It doessss

OpenStudy (whpalmer4):

Are there any other possible solutions?

OpenStudy (anonymous):

No I don't think so ? Solve -8 +\[\sqrt{5\times-5}\]=-3

OpenStudy (anonymous):

This one I need help on

OpenStudy (whpalmer4):

Look at the graph. Move over to 3 on the x-axis. Move up to the curve. Go to the left until you meet the y-axis. Where are you on the y-axis?